OpenStudy (adilalvi):

If 2x + 2x + 2x = 12 then what is the value of x in 2x - 1 3

OpenStudy (anonymous):

2

OpenStudy (lynfran):

whats 2x+2x+2x=?

OpenStudy (lynfran):

2x+2x+2x=12 6x=12 x=12/6 x=2 so sub x=2 into 2x-13

OpenStudy (lynfran):

o forget about the last line..error i misread the question

OpenStudy (adilalvi):

2x-13=0 2x=13 x=13/2=6.5 am i right.

OpenStudy (lynfran):

correct

OpenStudy (anonymous):

Is it 2x - 13 or 2x-1= 3?